A particular medication dosage is 45.0 mg/kg of body weight. What is the mass in mg of the medication a child weighing 25.1 lb. should receive?


1
Expert's answer
2021-02-01T04:05:54-0500

At first, pounds should be turned into kilogrammes:

m=25.1*0.45359237=11.385168487(kg)~11.39(kg)

Dosage=45.0*11.39=512.6(mg)